Trigonometry

How to graph a function of the for f(x)+f(X)on the same curve like for eg : to find the minimum value of sinx+cosx i need to graph it. . . . .? ? ?? ? ??

SINGH SINGH
14 Years agoGrade

1 Answer

Jitender Singh
12 Years ago
Ans: -2^1/2
Sol:
f(x) = sin x + cos x
sin (a + b) = sin a.cos b + cos a.sin b
sin x + cos x =2^1/2 sin (x +π/4)
f(x) =2^1/2 sin (x +π/4)
Now , you know the graph of sinx. Shift the graph from x = 0 to x = -π/4 and increase the amplitude to 2^1/2.
Good Luck!
Thanks & Regards
Jitender Singh
IIT Delhi
askIITians Faculty